OpenStudy (anonymous):

log[4]x=log[2]3 find x.

OpenStudy (anonymous):

Did you mean log (4^x) = log (2^3)? The answer to that is x=3/2

OpenStudy (anonymous):

i think he means log to base 4 of x and log to base 2 of 3.

OpenStudy (anonymous):

Oh thanks! Then, x=9